OpenStudy (anonymous):

im so frustrated please help me someone ! Find the ratio of the shortest side to the longest side . Write each ratio in lowest terms with whole numbers in the numerator and denominator. 1 1/2 ft, 2 1/4 ft, 2 ft

jimthompson5910 (jim_thompson5910):

what is the shortest side?

OpenStudy (anonymous):

1 1/2 ft maybe?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what is the longest side?

OpenStudy (anonymous):

2 1/4 ?

jimthompson5910 (jim_thompson5910):

so you divide 1 1/2 by 2 1/4

jimthompson5910 (jim_thompson5910):

and reduce fully

jimthompson5910 (jim_thompson5910):

1 1/2 = 1 + 1/2 2 1/4 = 2 + 1/4 Divide the two numbers